a body from height h strikes ground with 2 km/h  same weighted body from same height thrown down with initial velocity of 4km/h . Final v of 2 and body when striking ground??

Grade:12

1 Answers

Vikas TU
7 years ago
With 2km/hr.
2^2 = u^2 + 2gh
u = root(4 – 2gh)
 
With 4km/hr as initial velocty we get,
v^2 = 16 + 2gh
or
v = root(16+2gh)